# Prototype-Guided Diffusion: Visual Conditioning without External Memory

### v4

@@ Abstract
Diffusion models have emerged as a leading framework for high-quality image generation, offering stable training and strong performance across diverse domains. However, they remain computationally intensive, particularly during the iterative denoising process. Latent-space models like Stable Diffusion alleviate some of this cost by operating in compressed representations, though at the expense of fine-grained detail. More recent approaches such as Retrieval-Augmented Diffusion Models (RDM) address efficiency by conditioning denoising on similar examples retrieved from large external memory banks. While effective, these methods introduce drawbacks: they require costly storage and retrieval infrastructure, depend on static vision-language models like CLIP for similarity, and lack adaptability during training. We propose the Prototype Diffusion Model (PDM), a method that integrates prototype learning directly into the diffusion process for efficient and adaptive visual conditioning—without external memory. Instead of retrieving reference samples, PDM constructs a dynamic set of compact visual prototypes from clean image features using contrastive learning. These prototypes guide the denoising steps by aligning noisy representations with semantically relevant visual patterns, enabling efficient generation with strong semantic grounding. Experiments show that PDM maintains high generation quality while reducing computational and storage overhead, offering a scalable alternative to retrieval-based conditioning in diffusion models. Our code is publicly available at https://github.com/b-faye/pdm.
@@ I Introduction
Diffusion models have rapidly emerged as a dominant class of generative models, known for their high sample quality, training stability, and flexibility across diverse tasks. Since the introduction of Denoising Diffusion Probabilistic Models (DDPMs) [1], subsequent works have demonstrated their superiority over GAN-based approaches [2, 3]. These models have since been extended to tasks such as unconditional generation [4], super-resolution [5], inpainting [6], and notably, text-to-image generation [7, 8].
Despite their effectiveness, diffusion models are notoriously slow at both training and inference, owing to the iterative nature of the denoising process. To mitigate this, models such as Stable Diffusion [8] propose operating in a learned latent space, significantly reducing computational requirements. However, this efficiency comes at the cost of fine spatial details and often requires powerful encoders and decoders to preserve semantic fidelity.
More recently, methods such as Retrieval-Augmented Diffusion Models (RDM) [9] aim to condition the denoising process using external visual memories. By retrieving semantically similar image patches or examples via pretrained vision-language models (e.g., CLIP [10]), RDM dynamically steers generation toward more contextually relevant outputs. While this improves controllability and speeds up convergence, it introduces dependencies on large-scale memory banks and fixed feature extractors. These limitations raise concerns regarding memory cost, system latency, and adaptability during training.
An orthogonal line of work—ProtoDiffusion [11]—attempts to incorporate prototype learning into diffusion training by precomputing class-specific prototypes and using them as conditioning embeddings. Though effective in accelerating convergence, this two-stage approach decouples prototype learning from diffusion training, leading to rigidity in prototype representations and a lack of adaptability to evolving feature distributions. Moreover, class-level conditioning limits scalability in fine-grained or multimodal settings such as text-to-image generation.
We introduce the Prototype Diffusion Model (PDM), a unified framework that integrates online prototype learning into the diffusion training loop for efficient, context-aware generation—without relying on external memory or fixed embeddings. In contrast to RDM, our model dynamically constructs prototypes from clean intermediate features using contrastive objectives, and reuses them as visual anchors during the noisy denoising steps. Unlike ProtoDiffusion, our prototypes are learned jointly and continuously, adapting to the evolving feature space throughout training. This tightly-coupled prototype-guided mechanism enhances semantic grounding while maintaining efficiency. Our contributions are:
  • We propose a novel prototype-guided conditioning mechanism for diffusion models that avoids reliance on external retrieval systems or pretrained vision-language encoders.
  • Our method enables joint prototype learning and generation, allowing prototypes to adapt dynamically during training rather than relying on static pretraining.
  • We demonstrate that PDM matches or exceeds the generation quality of retrieval-augmented methods while significantly reducing memory overhead and inference latency.
  • We validate our approach on conditional image generation benchmarks, showing consistent gains in quality and speed compared to both retrieval-based and prototype-initialized baselines.
@@ II-A Diffusion Models
Diffusion models have emerged as a leading paradigm in generative modeling, delivering strong performance in tasks such as image generation, inpainting, super-resolution, and text-to-image synthesis. Following the foundational Denoising Diffusion Probabilistic Models (DDPMs) by Ho et al. [1], numerous advances have improved their training stability and sample quality. Latent Diffusion Models (LDMs) [8] reduce the computational load by operating in compressed latent spaces, enabling high-resolution synthesis with fewer resources. Further refinements such as consistency models [12] and SDXL [13] have demonstrated the scalability and expressiveness of diffusion-based generation. Nonetheless, the denoising process remains inherently slow due to the requirement for hundreds of sequential steps, making inference expensive. While latent-space methods partially alleviate this, they often compromise spatial detail and semantic fidelity. To address these shortcomings, researchers have proposed leveraging external context to augment and guide the denoising trajectory.
@@ II-B Retrieval-Augmented Diffusion Models
Retrieval-augmented approaches aim to improve generation quality and speed by incorporating semantically relevant information retrieved from external memory banks. Blattmann et al. [9] introduced Retrieval-Augmented Diffusion Models (RDMs), which retrieve contextually similar image patches to guide the diffusion model using CLIP-based feature alignment. Similarly, IRDiff [14] demonstrates retrieval-augmented molecular generation by retrieving known protein-ligand interactions, improving data efficiency and generalization. These methods enable controllable generation and reduce the burden on the denoiser by introducing external visual priors. However, they introduce new challenges: dependence on large, static memory systems, high retrieval latency, and limited adaptability to evolving feature distributions. Moreover, reliance on frozen pretrained encoders (e.g., CLIP [10]) constrains the flexibility of conditioning and may cause semantic mismatches between the prompt and retrieved examples. These limitations motivate internal memory mechanisms that can offer similar semantic grounding without relying on costly retrieval infrastructure.
@@ II-C Prototype Learning in Diffusion Models
Prototype learning provides a compact and interpretable alternative for guiding generative models without external databases. Baykal et al. [11] proposed ProtoDiffusion, incorporating class-specific prototypes as conditioning inputs to accelerate convergence and improve semantic alignment. However, their two-stage framework learns prototypes independently from the diffusion process, resulting in rigid, static prototypes that may fail to capture evolving feature distributions during training. In few-shot and continual learning contexts, Du et al. [15] and Doan et al. [16] introduced task-adaptive prototypes and class-conditioned diffusion to mitigate domain shifts and forgetting. While these approaches support dynamic conditioning, they either focus on discriminative tasks or rely on coarse, class-level prototypes—limiting fine-grained generation control and adaptability. Crucially, existing methods do not fully integrate prototype learning within the diffusion process itself.
Our proposed Prototype Diffusion Model (PDM) addresses this gap by introducing a fully joint training framework where patch-level prototypes are learned online from clean features and used to guide noisy denoising steps through a contrastive alignment objective. This enables PDM to maintain semantic consistency, support fine-grained conditioning, and reduce reliance on external memory—all while adapting prototypes dynamically throughout training.
@@ III-A Diffusion Models
Diffusion models are a class of generative models that learn to synthesize data by inverting a stochastic noising process. A prominent and widely used variant is the Denoising Diffusion Probabilistic Model (DDPM) [17, 1], which formulates generation as the reversal of a predefined forward diffusion process.
@@ III-B Retrieval-Augmented Diffusion Models
Retrieval-Augmented Diffusion Models (RDMs) enhance classical diffusion models by incorporating a retrieval mechanism that conditions generation on external, real examples. This semi-parametric approach supplements the learned generative model with non-parametric memory, enabling improved sample diversity, controllability, and generalization [8, 9, 18].
Given an input image $x$, its latent representation $z=E(x)$ is computed using a pretrained encoder $E$, typically from an autoencoder such as VQ-GAN [19]. A forward diffusion process is then applied in the latent space to produce noisy samples $z_{1},\dots,z_{T}$, analogous to the DDPM framework.
To enrich the generative process with semantic context, a retrieval function $\xi_{k}(x,D)$ selects the $k$ nearest neighbors of $x$ from a dataset $D$, based on similarity in a learned embedding space. This embedding space is induced by a pre-trained model $\phi$, such as CLIP [10], and the retrieval is defined as:
@@
where conditioning is typically implemented using cross-attention layers that attend to the retrieved features during noise prediction [8].
@@ III-C Prototype Learning in Diffusion Models
Prototype-based learning has become a cornerstone of few-shot learning, where the objective is to generalize to novel classes from only a few labeled examples. One of the most influential approaches in this domain is the Prototypical Network (ProtoNet) [20], which embeds both support and query samples into a common metric space and performs classification based on distances to class prototypes.
@@
Let $z_{c}$ denote the prototype for the target class $c$, and let $x_{t}$ be a noisy sample at timestep $t$. The denoising model can be conditioned on $z_{c}$ to produce a prototype-guided noise estimate $\epsilon_{\theta}(x_{t},t,z_{c})$. Following the classifier-free guidance strategy [21], the final guided noise estimate is computed as:
@@ IV Method
The second component of PDM is the denoising network $f_{\theta}$, implemented as a U-Net [22], which takes as input a noisy image and is conditioned on the selected prototype. To incorporate the prototype into the U-Net, cross-attention layers are added such that the prototype $e_{x}$ acts as both key and value, while the query originates from the output of the previous block processing the noisy image. This enables the network to guide the denoising process using the most relevant semantic concept.
@@ V Experiments
We conduct our experiments on CIFAR-10 [23], STL-10 [24], EuroSAT [25], and Tiny ImageNet [26] datasets using 2 NVIDIA A100 GPUs. Table I summarizes the key characteristics of the datasets, including the number of classes, total number of images, image resolution, and train/test splits.
Dataset # Classes # Training Images # Test Images Resolution
CIFAR-10 10 50,000 10,000 $32\times 32$
STL-10 10 5,000 8,000 $96\times 96$
EuroSAT 10 21,600 5,400 $64\times 64$
Tiny ImageNet 200 100,000 10,000 $64\times 64$
TABLE I: Summary of datasets used in our experiments.
For the denoiser (U-Net), we use the same compact architecture across all experiments, with a total block configuration of [128, 256, 256, 256] channels for successive downsampling and upsampling stages. Downsampling progressively increases the number of channels while reducing spatial resolution, whereas upsampling mirrors this process to recover the original resolution.
We incorporate the prototypes at the bottleneck, immediately after the final downsampling stage, using an attention mechanism. This design choice is motivated by the fact that the bottleneck captures the highest-level semantic features with the smallest spatial footprint, making it computationally efficient to attend to prototypes while maximizing their influence on global structure and class-specific semantics. By conditioning the most abstract feature representation on learned prototypes, we guide the generative process toward both semantic fidelity and intra-class consistency without imposing significant computational overhead.
We compare PDM and s-PDM against two relevant baselines. The first is DDPM, a standard U-Net–based diffusion model without prototypes, which can be viewed as an ablation of PDM where the prototype mechanism is removed. This comparison highlights that removing prototypes limits performance and that prototype-based conditioning can bring significant improvements. The second is ProtoDiffusion, a two-stage method in which prototypes are learned separately, frozen, and then injected into the bottleneck of the U-Net denoiser. While this approach incorporates prototypes, its separate training and freezing strategy make it computationally costly and unable to adapt prototypes to distribution shifts during denoiser training, which can progressively reduce their relevance. Our encoder $f_{\phi}$ uses a lightweight CNN with four convolutional layers followed by adaptive average pooling to extract features suitable for prototype learning.
For evaluation, we use three widely adopted generative quality metrics: Inception Score (IS) [27], Fréchet Inception Distance (FID) [25], and Kernel Inception Distance (KID) [28]. The IS measures both the diversity and semantic clarity of generated images by computing the KL divergence between the conditional label distribution predicted by a pretrained Inception network and its marginal distribution; it does not require real target images but relies on the classifier’s learned label space. In contrast, FID compares the distribution of real and generated images in the feature space of the Inception network by modeling both as multivariate Gaussians; it explicitly requires real target samples and evaluates how close the generated distribution is to the real one. Similarly, KID compares real and generated images in the Inception feature space but uses the squared Maximum Mean Discrepancy (MMD) with a polynomial kernel instead of Gaussian assumptions; like FID, it requires real target samples, but unlike FID, it is unbiased for a finite number of samples.
V-A Comparison of PDM, s-PDM, and Baseline Models on Generative Quality Metrics
@@
PDM (ours) 8.35 8.10 0.007 10.28 21.30 0.011 4.40 13.50 0.008 11.20 25.40 0.03
s-PDM (ours) 8.20 6.58 0.004 11.77 20.32 0.007 4.52 11.29 0.005 11.65 23.00 0.02
TABLE II: Quantitative comparison of generative quality across datasets using Inception Score (IS, $\uparrow$), Fréchet Inception Distance (FID, $\downarrow$), and Kernel Inception Distance (KID, $\downarrow$). For all datasets, the number of prototypes used equals the number of classes. Best results per column are in bold.
Table II shows a comprehensive comparison of generative performance across four datasets using IS, FID, and KID. Our proposed methods, PDM and s-PDM, consistently outperform the baselines DDPM and ProtoDiffusion on most metrics, demonstrating the effectiveness of incorporating prototypes within the diffusion framework.
In particular, s-PDM achieves the best FID and KID scores on all datasets, indicating superior generation quality and distributional alignment with the target data. This improvement can be attributed to the supervised nature of s-PDM: each class is associated with a fixed prototype that directly guides the generation process. As FID and KID explicitly measure the similarity between generated samples and real data distributions, having class-specific prototypes ensures more relevant and accurate sample generation.
On the other hand, IS does not explicitly consider class labels during evaluation, which explains why s-PDM’s IS is sometimes slightly lower than PDM or ProtoDiffusion. PDM, being unsupervised, optimizes to cluster similar images without label guidance, resulting in slightly better IS values that reflect sample diversity but less precise class alignment.
This difference is visually supported by Figure 1, where we apply PCA on the output features of $f_{\phi}$ and color-code samples by CIFAR-10 classes. The s-PDM clusters show clearer class separation than PDM, highlighting the advantage of supervised prototype learning in structuring the latent space according to semantic categories. In contrast, PDM’s parameters are randomly initialized and updated without class supervision, focusing solely on grouping similar images regardless of class labels.
Overall, these results confirm that supervised prototype guidance in s-PDM leads to more semantically meaningful and distributionally faithful generation, improving key metrics sensitive to sample quality and fidelity
Figure 1: PCA visualization of the output features of $f_{\phi}$ colored by CIFAR-10 classes, showing cluster separation for PDM and s-PDM.
V-B Ablation Study: Number of Prototypes in PDM
To evaluate the sensitivity of PDM to the number of prototypes, we vary this hyperparameter while keeping all other settings fixed. Since CIFAR-10 has $10$ classes, the default configuration uses $10$ prototypes (one per class). We investigate smaller values ($3,5,7$) and larger values ($13,15,20$) to observe how performance changes when the prototype count deviates from the number of classes.
# Prototypes IS $\uparrow$ FID $\downarrow$ KID $\downarrow$
@@ V-B Ablation Study: Number of Prototypes in PDM
TABLE III: Effect of varying the number of prototypes in PDM on CIFAR-10. Performance is measured using Inception Score (IS), Fréchet Inception Distance (FID), and Kernel Inception Distance (KID).
Table III shows that reducing the number of prototypes below the number of classes leads to a clear drop in performance across all metrics. For example, decreasing from $10$ to 33 prototypes reduces IS from $8.35$ to $7.50$ and increases FID from $8.10$ to $11.90$, indicating lower generation quality and poorer alignment with the real data distribution. This degradation occurs because fewer prototypes cannot represent the semantic diversity of the dataset, forcing multiple classes to share a prototype and thereby merging distinct modes in the data distribution.
When increasing the number of prototypes beyond $10$, IS improves slightly up to $15$ prototypes, reflecting a modest gain in sample diversity. However, FID and KID also decrease slightly, suggesting better fidelity and distributional alignment. At 2020 prototypes, IS drops again, likely due to over-fragmentation of the latent space: too many prototypes may split coherent clusters, leading to redundant or less meaningful centroids without adding new semantic structure.
These results are highly relevant because they demonstrate the importance of matching the number of prototypes to the inherent class structure of the dataset. The optimal setting ($10$ prototypes for CIFAR-10) aligns naturally with the number of classes, balancing diversity (high IS) with fidelity (low FID and KID). This finding supports the idea that prototypes in PDM should reflect the dataset’s underlying semantic granularity rather than being arbitrarily chosen.
V-C Qualitative Results: Sample Generations
(a) CIFAR-10
(b) EuroSAT
(c) STL-10
(d) PDM
(e) CIFAR-10
(f) EuroSAT
(g) STL-10
(h) s-PDM
Figure 3: Qualitative comparison of generated samples from PDM (top) and s-PDM (bottom) across CIFAR-10, EuroSAT, and STL-10. Each subfigure displays an $4\times 4$ grid of randomly generated images obtained after $500$ training epochs using $T=1000$ denoising steps. While s-PDM produces slightly sharper and more class-aligned samples, PDM still achieves visually plausible generations without supervision.
After training for $500$ epochs, we generate $4\times 4$ image grids by sampling from the model using Algorithm 2 with $T=1000$ denoising steps. For each generation, prototypes are selected randomly from the learned set. Figure 3 presents qualitative examples on CIFAR-10, EuroSAT and STL-10 for both PDM (unsupervised) and s-PDM (supervised).
Despite s-PDM achieving superior quantitative scores (Tables II), the visual fidelity of PDM remains competitive. The unsupervised variant captures coherent structures, realistic textures, and reasonable semantic consistency, even without explicit label guidance. This observation suggests that, in many practical scenarios where labeled data is limited or unavailable, an unsupervised prototype diffusion approach can still yield perceptually compelling generations while maintaining reduced training complexity.

### v5

@@ Abstract
Diffusion models achieve state-of-the-art image generation but remain computationally costly due to iterative denoising. Latent-space models like Stable Diffusion reduce overhead yet lose fine detail, while retrieval-augmented methods improve efficiency but rely on large memory banks, static similarity models, and rigid infrastructures.
We introduce the Prototype Diffusion Model (PDM), which embeds prototype learning into the diffusion process to provide adaptive, memory-free conditioning. Instead of retrieving references, PDM learns compact visual prototypes from clean features via contrastive learning, then aligns noisy representations with semantically relevant patterns during denoising. Experiments demonstrate that PDM sustains high generation quality while lowering computational and storage costs, offering a scalable alternative to retrieval-based conditioning.
@@ I Introduction
Diffusion models have rapidly become a dominant generative framework, known for high sample quality, training stability, and flexibility across tasks. Since the introduction of Denoising Diffusion Probabilistic Models (DDPMs) [ho2020denoising], they have surpassed GAN-based approaches [maze2023diffusion, dhariwal2021diffusion] and been extended to unconditional generation [song2020denoising], super-resolution [saharia2022photorealistic], inpainting [lugmayr2022repaint], and text-to-image generation [zhou2023shifted, rombach2022high].
Despite these advances, diffusion models remain slow due to iterative denoising. Latent-space methods such as Stable Diffusion [rombach2022high] mitigate this cost but lose spatial detail and require powerful encoders/decoders. Retrieval-augmented approaches (RDM) [blattmann2022retrieval] improve efficiency by conditioning on examples retrieved via pretrained vision-language models (e.g., CLIP [radford2021learning]), but incur large memory/storage costs and rely on fixed features, limiting adaptability. An orthogonal direction, ProtoDiffusion [baykal2024protodiffusion], introduces prototype-based conditioning but in a two-stage pipeline: class-specific prototypes are precomputed and remain static, hindering adaptability and scalability to fine-grained or multimodal settings.
We address these limitations with the Prototype Diffusion Model (PDM), which integrates online prototype learning directly into the diffusion process. Unlike ProtoDiffusion, PDM requires no labels and is fully unsupervised: prototypes are dynamically constructed from clean intermediate features using contrastive objectives, then reused as semantic anchors during noisy denoising steps. This joint, continuous learning allows prototypes to adapt alongside evolving representations, improving semantic grounding and efficiency without external memory. Furthermore, when labels are available, we extend PDM to a supervised variant, s-PDM, where class-specific prototypes are fixed by labels, eliminating prototype selection and compactness loss. Our contributions are:
  • We propose PDM, the first diffusion model that integrates online, unsupervised prototype learning into the denoising process, enabling efficient and adaptive conditioning without external memory or annotations.
  • We extend PDM to a supervised setting with s-PDM, where label-driven prototypes further simplify training and remove the need for compactness regularization.
  • Through extensive experiments, we show that both PDM and s-PDM outperform standard DDPMs (no prototypes) and ProtoDiffusion (two-stage static prototypes), validating their scalability and effectiveness.
@@ II-A Diffusion Models
Diffusion models have become a dominant generative framework, offering strong performance in image synthesis, inpainting, and super-resolution since DDPMs [ho2020denoising]. Latent diffusion [rombach2022high] improves efficiency by operating in compressed spaces, while recent advances such as consistency models [song2023] and SDXL [podell2024] demonstrate scalability. However, the denoising process remains slow due to many sequential steps, and latent-space acceleration often sacrifices spatial fidelity. These limitations motivate methods that incorporate additional context to guide denoising.
@@ II-B Retrieval-Augmented Diffusion Models
Retrieval-based approaches improve quality and controllability by conditioning diffusion on externally retrieved examples. RDM [blattmann2022retrieval] aligns noisy features with CLIP-based retrieved patches, and IRDiff [huang2024interaction] applies similar ideas to molecular generation. While effective, these methods rely on large static memory banks, frozen encoders [radford2021learning], and incur retrieval latency, reducing adaptability to evolving representations. Such constraints motivate lighter, internal conditioning mechanisms.
@@ II-C Prototype Learning in Diffusion Models
Prototype-based conditioning offers compact alternatives to external memory. ProtoDiffusion [baykal2024protodiffusion] employs class-specific prototypes but learns them in a separate, static stage, limiting adaptability. Other works use prototypes for few-shot or continual learning [du2024protodiff, doan2023class], yet typically rely on coarse class-level structures or focus on discriminative tasks. Existing methods thus lack fully joint, fine-grained prototype learning integrated into diffusion.
Our Prototype Diffusion Model (PDM) addresses this gap by learning patch-level prototypes online from clean features and aligning them with noisy representations during denoising, enabling adaptive conditioning without external memory.
@@ III-A Diffusion Models
Diffusion models are a class of generative models that learn to synthesize data by inverting a stochastic noising process. A prominent and widely used variant is the Denoising Diffusion Probabilistic Model (DDPM) [sohl2015deep, ho2020denoising], which formulates generation as the reversal of a predefined forward diffusion process.
@@ III-B Retrieval-Augmented Diffusion Models
Retrieval-Augmented Diffusion Models (RDMs) enhance classical diffusion models by incorporating a retrieval mechanism that conditions generation on external, real examples. This semi-parametric approach supplements the learned generative model with non-parametric memory, enabling improved sample diversity, controllability, and generalization [rombach2022high, blattmann2022retrieval, hu2023reveal].
Given an input image $x$, its latent representation $z=E(x)$ is computed using a pretrained encoder $E$, typically from an autoencoder such as VQ-GAN [esser2021taming]. A forward diffusion process is then applied in the latent space to produce noisy samples $z_{1},\dots,z_{T}$, analogous to the DDPM framework.
To enrich the generative process with semantic context, a retrieval function $\xi_{k}(x,D)$ selects the $k$ nearest neighbors of $x$ from a dataset $D$, based on similarity in a learned embedding space. This embedding space is induced by a pre-trained model $\phi$, such as CLIP [radford2021learning], and the retrieval is defined as:
@@
where conditioning is typically implemented using cross-attention layers that attend to the retrieved features during noise prediction [rombach2022high].
@@ III-C Prototype Learning in Diffusion Models
Prototype-based learning has become a cornerstone of few-shot learning, where the objective is to generalize to novel classes from only a few labeled examples. One of the most influential approaches in this domain is the Prototypical Network (ProtoNet) [snell2017prototypical], which embeds both support and query samples into a common metric space and performs classification based on distances to class prototypes.
@@
Let $z_{c}$ denote the prototype for the target class $c$, and let $x_{t}$ be a noisy sample at timestep $t$. The denoising model can be conditioned on $z_{c}$ to produce a prototype-guided noise estimate $\epsilon_{\theta}(x_{t},t,z_{c})$. Following the classifier-free guidance strategy [ho2022classifier], the final guided noise estimate is computed as:
@@ IV Method
The second component of PDM is the denoising network $f_{\theta}$, implemented as a U-Net [ronneberger2015u], which takes as input a noisy image and is conditioned on the selected prototype. To incorporate the prototype into the U-Net, cross-attention layers are added such that the prototype $e_{x}$ acts as both key and value, while the query originates from the output of the previous block processing the noisy image. This enables the network to guide the denoising process using the most relevant semantic concept.
@@ V Experiments
We evaluate PDM and its supervised variant s-PDM on CIFAR-10 [krizhevsky2009learning], STL-10 [coates2011analysis], EuroSAT [helber2019eurosat], and Tiny ImageNet [le2015tiny] using two NVIDIA A100 GPUs.
V-A Experimental Setup
All experiments use a compact U-Net with channel configuration [128, 256, 256, 256]. Prototypes are injected at the bottleneck through a single cross-attention block, enabling global semantic conditioning with minimal computational cost. The feature extractor $f_{\phi}$ is a lightweight 4-layer CNN with adaptive average pooling.
We compare our models to two baselines: (i) a standard DDPM (no prototypes) and (ii) ProtoDiffusion, where prototypes are learned in a separate stage and kept frozen. These comparisons allow us to isolate the effect of joint prototype learning during diffusion, which is the primary novelty of PDM. Importantly, our goal is not to surpass state-of-the-art diffusion performance, but to demonstrate that learning prototypes jointly with the denoiser leads to systematically improved generative quality.
Performance is reported using Inception Score (IS) [salimans2016improved], Fréchet Inception Distance (FID), and Kernel Inception Distance (KID) [binkowski2018demystifying].
V-B Quantitative Evaluation
@@
PDM (unsup.) 8.35 8.10 0.007 10.28 21.30 0.011 4.40 13.50 0.008 11.20 25.40 0.03
s-PDM (sup.) 8.20 6.58 0.004 11.77 20.32 0.007 4.52 11.29 0.005 11.65 23.00 0.02
TABLE I: Comparison of DDPM, ProtoDiffusion, PDM, and s-PDM across multiple datasets. Best values are in bold.
Table I shows that both PDM and s-PDM outperform the baselines across all datasets and metrics. The gains are especially clear for FID and KID, which require matching the real data distribution.
Key finding: jointly learning prototypes within the diffusion process (PDM) consistently improves generative fidelity over DDPM and ProtoDiffusion. This confirms that prototypes must remain trainable and aligned with the evolving denoiser features; otherwise, as in ProtoDiffusion, they gradually lose relevance.
s-PDM obtains the strongest results overall due to label-informed prototype assignment. Its advantage is most visible on FID/KID, where class-consistent conditioning directly improves distributional accuracy. In contrast, IS does not rely on real samples and favors diversity, which slightly benefits the unsupervised PDM.
Figure 1 illustrates this difference: PCA of $f_{\phi}$ features reveals sharper class separation in s-PDM, while PDM forms compact but unlabeled semantic groups. This confirms that supervised prototypes enforce stronger structure in latent space.
Figure 1: PCA of $f_{\phi}$ features on CIFAR-10. s-PDM shows clearer class separation, while PDM forms semantic clusters without labels.
V-C Effect of the Number of Prototypes
To assess the importance of prototype granularity, we vary the number of prototypes in PDM while keeping all other settings fixed.
Prototypes IS $\uparrow$ FID $\downarrow$ KID $\downarrow$
@@ V-C Effect of the Number of Prototypes
TABLE II: Ablation on the number of prototypes (CIFAR-10).
Results (Table II) show that using fewer prototypes than the dataset’s intrinsic semantic modes reduces performance, as multiple classes collapse into a single centroid. Increasing prototypes beyond the number of classes yields marginal gains but eventually causes over-fragmentation. These findings confirm that prototypes should match the dataset’s semantic granularity—a desirable property for adaptive or continual generative modeling.
V-D Visualization of Generated Samples
Figure 2 presents randomly generated $2\times 2$ image grids from four models trained on STL-10: DDPM, ProtoDiffusion, PDM, and s-PDM. All four models are shown side by side to emphasize differences in semantic coherence and prototype-guided generation.
(a) DDPM
(b) ProtoDiffusion
(c) PDM (unsupervised)
(d) s-PDM (supervised)
Figure 2: Random $2\times 2$ sample generations from models trained on STL-10 after 500 training epochs. Baseline models (DDPM and ProtoDiffusion) produce plausible but less semantically coherent images. Prototype-guided methods (PDM and s-PDM) demonstrate stronger semantic structure, with s-PDM further improving class consistency and visual fidelity. This highlights the impact of jointly learning prototypes on guiding the diffusion process.
These visualizations confirm that jointly learned prototypes substantially improve generation quality. While DDPM and ProtoDiffusion produce plausible images, they lack semantic consistency and class-specific alignment. PDM, even without labels, captures coherent structures and realistic textures, demonstrating that prototypes learned online can effectively guide the denoiser. s-PDM further enhances sharpness and class alignment by leveraging label-informed prototypes. Together, these results illustrate the central contribution of our work: learning prototypes jointly with the diffusion model provides strong semantic conditioning that improves both visual fidelity and structural coherence.